Consider a 1.0-L solution that is 0.830 M HNO2 and 0.37 M NaNO2 at 25 °C. What is the pH of this solution before and after 0.011 moles of LiOH have been added? Ka of HNO2 is 4.5×10-4.

initial pH: 3.7, final pH: 2.98.

initial pH: 3.35, final pH: 2.98.

initial pH: 3.0, final pH: 3.68.

initial pH: 3.7, final pH: 3.72.

initial pH: 3.0, final pH: 3.01.
Answer #1

1)

Lets calculate the initial pH

Ka = 4.5*10^-4

pKa = - log (Ka)

= - log(4.5*10^-4)

= 3.347

use:

pH = pKa + log {[conjugate base]/[acid]}

= 3.347+ log {0.37/0.83}

= 3.00

2)

Lets calculate the final pH

mol of LiOH added = 0.011 mol

HNO2 will react with OH- to form NO2-

Before Reaction:

mol of NO2- = 0.37 M *1.0 L

mol of NO2- = 0.37 mol

mol of HNO2 = 0.83 M *1.0 L

mol of HNO2 = 0.83 mol

after reaction,

mol of NO2- = mol present initially + mol added

mol of NO2- = (0.37 + 0.011) mol

mol of NO2- = 0.381 mol

mol of HNO2 = mol present initially - mol added

mol of HNO2 = (0.83 - 0.011) mol

mol of HNO2 = 0.819 mol

Ka = 4.5*10^-4

pKa = - log (Ka)

= - log(4.5*10^-4)

= 3.347

since volume is both in numerator and denominator, we can use mol instead of concentration

use:

pH = pKa + log {[conjugate base]/[acid]}

= 3.347+ log {0.381/0.819}

= 3.01

Answer:

initial pH: 3.0, final pH: 3.01.